OpenStudy (anonymous):

6. There are 20 boys and 20 girls in a class. The average percentage marks of the students in the class are more than 70, but less than 85 for a particular monthly test. The average percentage marks scored by the boys is more than 60, but less than 75. What is the range of the average percentage marks scored by the girls? A. 65≤y≤110 B. 140≤y≤170 C. 60≤y≤75 D. 75≤y≤110

OpenStudy (anonymous):

C.

OpenStudy (anonymous):

there are as many boys as girls so the percentage should be the same as the boys

OpenStudy (anonymous):

It looks like it would have to be D, since the average is between 70 and less than 85, however the boys barely make the 75 mark, so to offset that, the girls would have to be higher to meet the mean of the problem.

OpenStudy (anonymous):

wrong answer, @sunsetlove and good rationale @jessicajuanng. The total score of the class will be TOT=40*AVG_PCT_CLASS, right? The Boys' contribution is 20*AVG_PCT_BOYS And the Girls': 20*AVG_PCT_GIRLS Boys' and Gals' contribution must be the class total, then: AVG_PCT_GIRLS=(40*AVG_PCT_CLASS-20*AVG_PCT_BOYS)/20=2*AVG_PCT_CLASS-AVG_PCT_BOYS With this, the lowest percentage for girls' will be when the total of the class is lowest and the boys' is highest The other way around for the girls' highest percentage Now you get back to me with the solution